# From Word Vectors to Multimodal Embeddings: Techniques, Applications, and Future Directions For Large Language Models

### v1

@@ V-D1 Computational Cost of Training Large Embedding Matrices
The increasing scale of data used in natural language processing (NLP) has led to a proportional increase in embedding matrix sizes, creating significant computational challenges. This is particularly evident in pre-trained language models like BERT and GPT, which often contain billions of parameters. The memory footprint of these models is substantial, posing difficulties for deployment on resource-constrained devices and impacting both training and inference times. For instance, [13] emphasizes the need for efficient language representations and improved evaluation methods, while [77] focuses on compressing large models like BERT to reduce computational demands.
@@ V-D2 Efficient Training Techniques
Several techniques aim to reduce the computational cost of training large embedding matrices while preserving the quality of the learned embeddings [7]. These methods focus on optimizing the training process to achieve faster convergence and lower resource consumption.
@@
Noise-contrastive estimation (NCE) approximates the loss function in training word embeddings, significantly accelerating the process. [10] demonstrates how NCE enables high-quality word embedding training with reduced data and computational needs by simplifying loss computation, especially beneficial for large vocabularies and datasets.
@@
Subsampling reduces the number of training examples used in the embedding learning process, facilitating faster convergence and lower computational costs. [95] presents an efficient subsampling approach for neural word embeddings trained with hierarchical softmax, focusing on the most informative examples for quicker and potentially more accurate learning.
@@
Optimized batching strategies enhance training throughput by employing larger batch sizes, which is particularly useful for large language models. [96] discusses using large batches to improve embedding discriminativeness. Larger batch sizes lead to more stable parameter updates and faster convergence, although careful tuning is required to manage memory constraints and convergence stability.
@@
Parallel and distributed training techniques distribute computational workload across multiple GPUs or CPUs, drastically reducing training time for large models and datasets. [42] describes parallel training methods that enable scaling beyond previous model sizes, facilitating training for very large neural networks.
@@ V-D3 Hardware Acceleration for Embedding Training and Inference
Specialized hardware such as GPUs and TPUs offers significant acceleration for both training and inference of embedding models. These devices’ parallel processing capabilities are ideally suited for the matrix operations integral to embedding computations, expediting model experimentation and deployment in NLP applications [97].
@@ V-E1 Sources of Bias in Embedding Models
Biases in language models pose ethical challenges and can perpetuate societal inequalities. These biases arise from various sources, including the training data and the model architecture. Training data often reflects societal biases, as noted by [13]. If the corpus overrepresents specific demographics or viewpoints, the resulting embeddings may inherit and amplify these biases. Model architecture also contributes to bias; design choices, such as the objective function or the incorporation of context, may unintentionally favor certain types of information, leading to skewed representations. For instance, models that prioritize word co-occurrence patterns may inadvertently capture and reinforce biases present in the data [32].
@@ V-E2 Measuring and Quantifying Bias
Accurate measurement and quantification of bias are essential for understanding its nature and impact. Intrinsic and extrinsic evaluation metrics are commonly used. Intrinsic evaluations directly assess embeddings for bias, often through similarity measures or analogy tasks. For instance, [98] used word similarity tasks to compare translation-based and monolingual embeddings, revealing differences in how they capture ontological information. Extrinsic evaluations measure bias by examining its impact on downstream NLP tasks. Statistical tests, like the Word Embedding Association Test (WEAT), help quantify biased associations, while [99] used correlation analysis to explore the relationship between intrinsic and extrinsic evaluations. Analyzing bias’s differential impact on social groups is also essential; [91] investigated intersectional biases, examining how gender identity, sexual orientation, and social class intersect to influence biased representations.
@@ V-E3 Mitigating Bias
Bias mitigation in language models is an active area of research. Data augmentation can balance training datasets, minimizing the effects of skewed representations [15]. Debiasing techniques aim to neutralize biased associations within embedding spaces; for instance, [92] proposed a new metric for evaluating vector transformations. Additionally, adversarial training helps models become robust to biased inputs by making them insensitive to specific protected attributes, enhancing fairness in model predictions.
@@ V-E4 Ethical Implications of Biased Embeddings
Biased embeddings raise ethical concerns surrounding fairness, transparency, and accountability in NLP systems. Embeddings that perpetuate stereotypes or discriminate against certain groups can lead to harmful societal consequences. For example, [47] examined word embeddings for hate speech detection, emphasizing the importance of capturing context for accurate identification. Transparency is essential to understand the origins of biases and their impact on model decisions, while accountability ensures that those developing and deploying NLP systems take responsibility for mitigating bias and addressing potential harms.
@@ V-F1 Transfer Learning: Using Pre-trained Embeddings for Downstream Tasks
Pre-trained word embeddings have become a cornerstone of NLP, providing rich semantic information derived from extensive unlabeled text. This transfer learning approach has proven effective across diverse NLP applications. For instance, [11] demonstrates that LSTM-RNN sentence embeddings enhance web document retrieval performance, while [100] shows how neural word embeddings improve translation model effectiveness. Beyond retrieval tasks, [32] analyzes properties encoded in sentence embeddings, offering insights into their capture of content, order, and length. Pre-trained embeddings also benefit other areas: [24] employs word embeddings to tackle vocabulary mismatch in information retrieval, yielding notable improvements, and [40] highlights the effectiveness of BERT embeddings in knowledge base question answering.
@@ V-F2 Domain Adaptation: Fine-tuning Embeddings on Domain-Specific Data
While generic pre-trained embeddings provide a strong baseline, fine-tuning on domain-specific data can further enhance their utility, particularly for specialized vocabularies and semantic nuances. [101] demonstrates that embeddings trained on engineering texts outperform general-purpose models in engineering-related tasks. In the clinical domain, [102] finds that contextual embeddings pre-trained on clinical data achieve state-of-the-art performance in clinical concept extraction. Similarly, [37] explores the impact of domain-specific context on RNN language models, emphasizing the benefits of domain adaptation.
@@ V-F3 Cross-Lingual Transfer Learning: Leveraging Embeddings for Low-Resource Languages
Cross-lingual embeddings enable knowledge transfer from resource-rich to resource-poor languages by projecting words into a shared embedding space. [28] explores cross-lingual embeddings from bilingual lexicons to improve low-resource language models, while [30] introduces bilingual embeddings learned from unlabeled corpora and machine translation alignments, enhancing translation performance. Even with limited data, [103] shows that embeddings from neural machine translation models outperform monolingual embeddings by capturing conceptual and lexical-syntactic similarities. [60] further proposes an unsupervised method for cross-lingual embeddings using multilingual language models, achieving high quality with limited monolingual data and across varying domains.
@@ V-F4 Adaptive Language Modeling: Conditioning Language Models on Different Contexts
Adaptive language modeling allows models to condition on specific contexts, such as style or sentiment, for more controlled text generation. [104] explores weighting schemes for ELMo embeddings to improve downstream performance in varied contexts. [105] introduces LM-Switch, a lightweight approach for adapting language models to specific conditions by manipulating word embeddings, enabling nuanced language generation.
@@
Zero-shot learning with embeddings enables language models to tackle new tasks without task-specific training data. [67] introduces the Visual Semantic Embedding Probe (VSEP) for evaluating contextualized word embeddings in zero-shot visual semantic understanding, while [106] argues for evaluating language model embeddings on their direct linguistic reasoning capabilities, showing potential for zero-shot common-sense reasoning tasks.
{forest}
[…]
and Research Gaps, root, [Embedding Model
Compression, level1, [Compression Techniques, level2, [Knowledge Distillation [52, 28, 77], level2] [Weight Tying [46, 78, 79], level2] [Quantization [80], level2] [Pruning [81], level2] [Subspace Embedding Methods
[82, 83, 84], level2] ] [Evaluating Compression Methods
[77, 82], level2] ] [Interpretability and
Explainability
of Embeddings, level1, [The Black Box Nature [55, 32], level2] [Methods for Understanding
Embeddings, level2, [Visualization Techniques [85, 86], level2] [Probing Tasks [87, 88, 55], level2] [Concept Activation Vectors
(CAVs) [89], level2] [Embedding Inversion Attacks [90], level2] ] [Using LLMs to Interpret Embeddings, level2] [Explainable Embeddings [91], level2] ] [Encoding
Numerical Information
in Language Models, level1, [Challenges of
Representing Numbers [92], level2] [Embedding Numerical Information, level2, [Default Embedding Methods [42], level2] [Exponential Embeddings [37], level2] [Floating-Point Embeddings [93], level2] ] [Specialized Architectures for
Numerical Reasoning [94], level2] [Evaluation of Numerical
Reasoning Capabilities [93], level2] ] [Efficient and Scalable
Training of
Embedding Models, level1, [Computational
Cost of Training [13, 77], level2] [Efficient Training
Techniques [10, 95, 96, 42], level2, [Noise-Contrastive Estimation [10], level2] [Subsampling Techniques [95], level2] [Optimized Batching Strategies [96], level2] [Parallel and Distributed
Training Methods [42], level2] ] [Hardware Acceleration
for Embedding Training
and Inference [97], level2] ] [Addressing
Bias and Ethical
Concerns in Embeddings, level1, [Sources of Bias in
Embedding Models [13, 32], level2] [Measuring and Quantifying Bias
[98, 99, 91], level2] [Mitigating Bias [15, 92], level2] [Ethical Implications
of Biased Embeddings [47], level2] ] [Adaptive
Language Modeling and
Transfer Learning
with Embeddings, level1, [Transfer Learning
[11, 100, 32, 24, 40], level2] [Domain Adaptation [101, 102, 37], level2] [Cross-Lingual Transfer Learning
[28, 30, 103, 60], level2] [Adaptive Language Modeling
[104, 105], level2] [Zero-Shot Learning with
Embeddings [67, 106], level2] ] [The Role of Embeddings
in Emerging Areas, level1, [Embodied AI [66, 107, 69], level2] [Cognitive Science
[74, 108, 109, 73, 75, 32], level2] [Reasoning and
Commonsense Knowledge
[110, 111, 26, 106, 43], level2] ] ]
@@ V-G1 Embodied AI: Grounding Language in Sensorimotor Experiences
Embodied AI seeks to ground language understanding in sensorimotor experiences, enabling agents to interact with and reason about the physical world. Multimodal embeddings are crucial for bridging linguistic instructions and robotic actions. [66] introduced an algorithm that learns a shared embedding space for point-cloud data, natural language, and manipulation trajectories, enhancing accuracy and inference in robotic tasks. This approach allows robots to generalize from past interactions with objects to manipulate new objects based on linguistic commands. Further advancing embodied AI, [107] proposed PaLM-E, a multimodal language model that integrates continuous sensor data (e.g., visual, state estimation) into a large language model, excelling in embodied tasks like robotic manipulation planning, visual question answering, and captioning. Kosmos-2 also contributes to this field by grounding text in visual information, such as bounding boxes, which supports complex multimodal tasks like phrase grounding and referring expression comprehension [69].
@@ V-G2 Cognitive Science: Modeling Human Language Processing with Embeddings
Embeddings have become a valuable tool in cognitive science for modeling human language processing and understanding the neural basis of language. Studies reveal a striking correspondence between brain activity and artificial embeddings. [74] observed that brain embeddings in Broca’s area share a similar geometry with contextual embeddings from deep language models, allowing cross-domain predictions and indicating a shared representational space. Further research, such as [109], reinforces this shared geometry’s predictive power, suggesting that embedding models provide a computational framework for language representation in the brain. Additionally, [73] explored subject-specific semantic representations, finding that language models can capture individual differences in processing personally familiar entities. [75] extended these insights by highlighting a correspondence between deep language model (DLM) layers and the temporal dynamics of brain activity during language processing. These findings suggest that artificial embeddings offer valuable insights into neural mechanisms underlying language processing [32].
@@
Integrating embeddings with probabilistic models and knowledge graphs enhances commonsense reasoning and knowledge representation in language models. [110] introduced a model-independent loss function that incorporates pre-trained language models into knowledge graph embeddings, improving knowledge graph completion accuracy. Additionally, [111] demonstrated that joint modeling of knowledge graphs and entity prediction with language models enhances performance on both tasks. [26] provided a comprehensive review of word embeddings and pre-trained language models, emphasizing their role in advancing NLP tasks. [106] proposed a paradigm that evaluates language models based on their capacity for direct linguistic reasoning, suggesting that embeddings function as implicit knowledge repositories. [26] also explored integrating pre-trained language models into sequence-to-sequence models, highlighting their effectiveness in tasks like machine translation and summarization. These studies illustrate the potential of combining embeddings with knowledge representation techniques to boost reasoning capabilities. Moreover, [43] introduced a generative model for word embeddings, providing a theoretical foundation for nonlinear models and explaining how embeddings support analogical reasoning. Fig 3 presents a taxonomy of advanced topics and research gaps covered in this section.

### v2

@@ V-D1 Computational Cost of Training Large Embedding Matrices
The increasing scale of data used in natural language processing (NLP) has led to a proportional increase in embedding matrix sizes, creating significant computational challenges. This is particularly evident in pre-trained language models like BERT and GPT, which often contain billions of parameters. The memory footprint of these models is substantial, posing difficulties for deployment on resource-constrained devices and impacting both training and inference times. For instance, [yong_intrinsic_2018] emphasizes the need for efficient language representations and improved evaluation methods, while [sanqiang_extreme_2019] focuses on compressing large models like BERT to reduce computational demands.
@@ V-D2 Efficient Training Techniques
Several techniques aim to reduce the computational cost of training large embedding matrices while preserving the quality of the learned embeddings [lebret_word_2016]. These methods focus on optimizing the training process to achieve faster convergence and lower resource consumption.
@@
Noise-contrastive estimation (NCE) approximates the loss function in training word embeddings, significantly accelerating the process. [a_learning_2013] demonstrates how NCE enables high-quality word embedding training with reduced data and computational needs by simplifying loss computation, especially beneficial for large vocabularies and datasets.
@@
Subsampling reduces the number of training examples used in the embedding learning process, facilitating faster convergence and lower computational costs. [stephan_training_2016] presents an efficient subsampling approach for neural word embeddings trained with hierarchical softmax, focusing on the most informative examples for quicker and potentially more accurate learning.
@@
Optimized batching strategies enhance training throughput by employing larger batch sizes, which is particularly useful for large language models. [chen_bge_2024] discusses using large batches to improve embedding discriminativeness. Larger batch sizes lead to more stable parameter updates and faster convergence, although careful tuning is required to manage memory constraints and convergence stability.
@@
Parallel and distributed training techniques distribute computational workload across multiple GPUs or CPUs, drastically reducing training time for large models and datasets. [andrew_modeling_2015] describes parallel training methods that enable scaling beyond previous model sizes, facilitating training for very large neural networks.
@@ V-D3 Hardware Acceleration for Embedding Training and Inference
Specialized hardware such as GPUs and TPUs offers significant acceleration for both training and inference of embedding models. These devices’ parallel processing capabilities are ideally suited for the matrix operations integral to embedding computations, expediting model experimentation and deployment in NLP applications [raha2023efficient].
@@ V-E1 Sources of Bias in Embedding Models
Biases in language models pose ethical challenges and can perpetuate societal inequalities. These biases arise from various sources, including the training data and the model architecture. Training data often reflects societal biases, as noted by [yong_intrinsic_2018]. If the corpus overrepresents specific demographics or viewpoints, the resulting embeddings may inherit and amplify these biases. Model architecture also contributes to bias; design choices, such as the objective function or the incorporation of context, may unintentionally favor certain types of information, leading to skewed representations. For instance, models that prioritize word co-occurrence patterns may inadvertently capture and reinforce biases present in the data [yossi_analysis_2017].
@@ V-E2 Measuring and Quantifying Bias
Accurate measurement and quantification of bias are essential for understanding its nature and impact. Intrinsic and extrinsic evaluation metrics are commonly used. Intrinsic evaluations directly assess embeddings for bias, often through similarity measures or analogy tasks. For instance, [felix_not_2014] used word similarity tasks to compare translation-based and monolingual embeddings, revealing differences in how they capture ontological information. Extrinsic evaluations measure bias by examining its impact on downstream NLP tasks. Statistical tests, like the Word Embedding Association Test (WEAT), help quantify biased associations, while [bin_evaluating_2019] used correlation analysis to explore the relationship between intrinsic and extrinsic evaluations. Analyzing bias’s differential impact on social groups is also essential; [shiva_evaluating_2023] investigated intersectional biases, examining how gender identity, sexual orientation, and social class intersect to influence biased representations.
@@ V-E3 Mitigating Bias
Bias mitigation in language models is an active area of research. Data augmentation can balance training datasets, minimizing the effects of skewed representations [matthias_context_2018]. Debiasing techniques aim to neutralize biased associations within embedding spaces; for instance, [t_language_2021] proposed a new metric for evaluating vector transformations. Additionally, adversarial training helps models become robust to biased inputs by making them insensitive to specific protected attributes, enhancing fairness in model predictions.
@@ V-E4 Ethical Implications of Biased Embeddings
Biased embeddings raise ethical concerns surrounding fairness, transparency, and accountability in NLP systems. Embeddings that perpetuate stereotypes or discriminate against certain groups can lead to harmful societal consequences. For example, [minni_comparison_2021] examined word embeddings for hate speech detection, emphasizing the importance of capturing context for accurate identification. Transparency is essential to understand the origins of biases and their impact on model decisions, while accountability ensures that those developing and deploying NLP systems take responsibility for mitigating bias and addressing potential harms.
@@ V-F1 Transfer Learning: Using Pre-trained Embeddings for Downstream Tasks
Pre-trained word embeddings have become a cornerstone of NLP, providing rich semantic information derived from extensive unlabeled text. This transfer learning approach has proven effective across diverse NLP applications. For instance, [hamid_deep_2015] demonstrates that LSTM-RNN sentence embeddings enhance web document retrieval performance, while [g_integrating_2015] shows how neural word embeddings improve translation model effectiveness. Beyond retrieval tasks, [yossi_analysis_2017] analyzes properties encoded in sentence embeddings, offering insights into their capture of content, order, and length. Pre-trained embeddings also benefit other areas: [debasis_word_2015] employs word embeddings to tackle vocabulary mismatch in information retrieval, yielding notable improvements, and [japa_question_2020] highlights the effectiveness of BERT embeddings in knowledge base question answering.
@@ V-F2 Domain Adaptation: Fine-tuning Embeddings on Domain-Specific Data
While generic pre-trained embeddings provide a strong baseline, fine-tuning on domain-specific data can further enhance their utility, particularly for specialized vocabularies and semantic nuances. [daniel_language_2021] demonstrates that embeddings trained on engineering texts outperform general-purpose models in engineering-related tasks. In the clinical domain, [yuqi_enhancing_2019] finds that contextual embeddings pre-trained on clinical data achieve state-of-the-art performance in clinical concept extraction. Similarly, [linda_effect_2019] explores the impact of domain-specific context on RNN language models, emphasizing the benefits of domain adaptation.
@@ V-F3 Cross-Lingual Transfer Learning: Leveraging Embeddings for Low-Resource Languages
Cross-lingual embeddings enable knowledge transfer from resource-rich to resource-poor languages by projecting words into a shared embedding space. [trevor_cross-lingual_2017] explores cross-lingual embeddings from bilingual lexicons to improve low-resource language models, while [will_bilingual_2013] introduces bilingual embeddings learned from unlabeled corpora and machine translation alignments, enhancing translation performance. Even with limited data, [felix_embedding_2014] shows that embeddings from neural machine translation models outperform monolingual embeddings by capturing conceptual and lexical-syntactic similarities. [takashi_unsupervised_2018] further proposes an unsupervised method for cross-lingual embeddings using multilingual language models, achieving high quality with limited monolingual data and across varying domains.
@@ V-F4 Adaptive Language Modeling: Conditioning Language Models on Different Contexts
Adaptive language modeling allows models to condition on specific contexts, such as style or sentiment, for more controlled text generation. [nils_alternative_2019] explores weighting schemes for ELMo embeddings to improve downstream performance in varied contexts. [chi_word_2023] introduces LM-Switch, a lightweight approach for adapting language models to specific conditions by manipulating word embeddings, enabling nuanced language generation.
@@
Zero-shot learning with embeddings enables language models to tackle new tasks without task-specific training data. [yue_language_2021] introduces the Visual Semantic Embedding Probe (VSEP) for evaluating contextualized word embeddings in zero-shot visual semantic understanding, while [nancy_rethinking_2020] argues for evaluating language model embeddings on their direct linguistic reasoning capabilities, showing potential for zero-shot common-sense reasoning tasks.
Advanced Topics and Research Gaps Embedding Model Compression Compression Techniques Knowledge Distillation [nils_making_2020, trevor_cross-lingual_2017, sanqiang_extreme_2019] Weight Tying [hakan_tying_2016, ofir_using_2016, kristina_how_2018] Quantization [tao2022compression] Pruning [wang2019structured] Subspace Embedding Methods [amit_lightweight_2023, xu_tensorgpt_2023, kartik_diverse_2014] Evaluating Compression Methods [sanqiang_extreme_2019, amit_lightweight_2023] Interpretability and Explainability of Embeddings The Black Box Nature [steven_analysing_2020, yossi_analysis_2017] Methods for Understanding Embeddings Visualization Techniques [dima_comparative_2018, rishabh_comparative_2021] Probing Tasks [alessio_contextual_2020, matthew_dissecting_2018, steven_analysing_2020] Concept Activation Vectors (CAVs) [tennenholtz_demystifying_2023] Embedding Inversion Attacks [li_sentence_2023] Using LLMs to Interpret Embeddings Explainable Embeddings [shiva_evaluating_2023] Encoding Numerical Information in Language Models Challenges of Representing Numbers [t_language_2021] Embedding Numerical Information Default Embedding Methods [andrew_modeling_2015] Exponential Embeddings [linda_effect_2019] Floating-Point Embeddings [xiaoxiao_floating-point_2024] Specialized Architectures for Numerical Reasoning [qiang_bilstm_2021] Evaluation of Numerical Reasoning Capabilities [xiaoxiao_floating-point_2024] Efficient and Scalable Training of Embedding Models Computational Cost of Training [yong_intrinsic_2018, sanqiang_extreme_2019] Efficient Training Techniques [a_learning_2013, stephan_training_2016, chen_bge_2024, andrew_modeling_2015] Noise-Contrastive Estimation [a_learning_2013] Subsampling Techniques [stephan_training_2016] Optimized Batching Strategies [chen_bge_2024] Parallel and Distributed Training Methods [andrew_modeling_2015] Hardware Acceleration for Embedding Training and Inference [raha2023efficient] Addressing Bias and Ethical Concerns in Embeddings Sources of Bias in Embedding Models [yong_intrinsic_2018, yossi_analysis_2017] Measuring and Quantifying Bias [felix_not_2014, bin_evaluating_2019, shiva_evaluating_2023] Mitigating Bias [matthias_context_2018, t_language_2021] Ethical Implications of Biased Embeddings [minni_comparison_2021] Adaptive Language Modeling and Transfer Learning with Embeddings Transfer Learning [hamid_deep_2015, g_integrating_2015, yossi_analysis_2017, debasis_word_2015, japa_question_2020] Domain Adaptation [daniel_language_2021, yuqi_enhancing_2019, linda_effect_2019] Cross-Lingual Transfer Learning [trevor_cross-lingual_2017, will_bilingual_2013, felix_embedding_2014, takashi_unsupervised_2018] Adaptive Language Modeling [nils_alternative_2019, chi_word_2023] Zero-Shot Learning with Embeddings [yue_language_2021, nancy_rethinking_2020] The Role of Embeddings in Emerging Areas Embodied AI [jaeyong_deep_2015, driess_palm-e_2023, peng_kosmos-2_2023] Cognitive Science [ariel_brain_2022, niu2024large, ariel_alignment_2024, andrea_family_2023, goldstein_correspondence_2022, yossi_analysis_2017] Reasoning and Commonsense Knowledge [mirza_language_2022, rajat_complementary_2020, mourad_word_2022, nancy_rethinking_2020, sanjeev_latent_2015]
@@ V-G1 Embodied AI: Grounding Language in Sensorimotor Experiences
Embodied AI seeks to ground language understanding in sensorimotor experiences, enabling agents to interact with and reason about the physical world. Multimodal embeddings are crucial for bridging linguistic instructions and robotic actions. [jaeyong_deep_2015] introduced an algorithm that learns a shared embedding space for point-cloud data, natural language, and manipulation trajectories, enhancing accuracy and inference in robotic tasks. This approach allows robots to generalize from past interactions with objects to manipulate new objects based on linguistic commands. Further advancing embodied AI, [driess_palm-e_2023] proposed PaLM-E, a multimodal language model that integrates continuous sensor data (e.g., visual, state estimation) into a large language model, excelling in embodied tasks like robotic manipulation planning, visual question answering, and captioning. Kosmos-2 also contributes to this field by grounding text in visual information, such as bounding boxes, which supports complex multimodal tasks like phrase grounding and referring expression comprehension [peng_kosmos-2_2023].
@@ V-G2 Cognitive Science: Modeling Human Language Processing with Embeddings
Embeddings have become a valuable tool in cognitive science for modeling human language processing and understanding the neural basis of language. Studies reveal a striking correspondence between brain activity and artificial embeddings. [ariel_brain_2022] observed that brain embeddings in Broca’s area share a similar geometry with contextual embeddings from deep language models, allowing cross-domain predictions and indicating a shared representational space. Further research, such as [ariel_alignment_2024], reinforces this shared geometry’s predictive power, suggesting that embedding models provide a computational framework for language representation in the brain. Additionally, [andrea_family_2023] explored subject-specific semantic representations, finding that language models can capture individual differences in processing personally familiar entities. [goldstein_correspondence_2022] extended these insights by highlighting a correspondence between deep language model (DLM) layers and the temporal dynamics of brain activity during language processing. These findings suggest that artificial embeddings offer valuable insights into neural mechanisms underlying language processing [yossi_analysis_2017].
@@
Integrating embeddings with probabilistic models and knowledge graphs enhances commonsense reasoning and knowledge representation in language models. [mirza_language_2022] introduced a model-independent loss function that incorporates pre-trained language models into knowledge graph embeddings, improving knowledge graph completion accuracy. Additionally, [rajat_complementary_2020] demonstrated that joint modeling of knowledge graphs and entity prediction with language models enhances performance on both tasks. [mourad_word_2022] provided a comprehensive review of word embeddings and pre-trained language models, emphasizing their role in advancing NLP tasks. [nancy_rethinking_2020] proposed a paradigm that evaluates language models based on their capacity for direct linguistic reasoning, suggesting that embeddings function as implicit knowledge repositories. [mourad_word_2022] also explored integrating pre-trained language models into sequence-to-sequence models, highlighting their effectiveness in tasks like machine translation and summarization. These studies illustrate the potential of combining embeddings with knowledge representation techniques to boost reasoning capabilities. Moreover, [sanjeev_latent_2015] introduced a generative model for word embeddings, providing a theoretical foundation for nonlinear models and explaining how embeddings support analogical reasoning. Fig 3 presents a taxonomy of advanced topics and research gaps covered in this section.